Compute turnaround time with preemptive priority scheduling. Also compute the average turnaround time. Show the Gantt chart and other details.

Process

CPU Burst time

Arrival time

Priority

P1

25

5

5

P2

10

10

3

P3

15

15

4

P4

20

20

2

P5

30

25

1
Answer #1

Total Turn Around Time = Completion Time - Arrival Time

Let highest number has highest priority. i.e. process with priority 5 has maximum priority and with 1 has lower priority.

So at t = 5

only P1 is there, it will start its execution and at t=10 P2 arrives but its priority is lesser than currenct executing process.

at t = 15 P3 arrives it also has lesser priority then P1.

at t=20 P4 arrives and t=25 P5 arrives. all has lower priority.

Burst time for P1 is 25 and it started at t=5 so it will end its execution at t=30.

So TAT for P1 = 30 -5 = 25

now P3 has priority 4 so it will start its execution at t=30,

completion time = 30+15 = 45

its arrival time = 15

so TAT for P3 = 45-15 =30

Now P2 start its execution at t=45

arrival time = 10

CPU burst time = 10

so completion time = 45 + 10 =55

TAT for P2 = 55-10 = 45

Now P4 start its execution at t=55

arrival time = 20

CPU burst time = 20

so completion time = 55 + 20 =75

TAT for P4 = 75-20 = 55

Now P5 start its execution at t=75

arrival time = 30

CPU burst time = 25

so completion time = 75 + 25 =100

TAT for P5 = 100-30 = 70




Hope this helps!

You can ask your doubts in comment section.

Kindly appreciate the help by upvoting the answer. Thank you!